Find P (12 < x < 18); µ = 12, σ = 5
Answer by ewatrrr(24785)

 
Hi
µ = 12, σ = 5
P (12 < x < 18) = P(z=(18-12)/5 - P(z =(12-12)/5 = P(z=1.2) -.5 = .8849 - .5 = .3849
Note:Using the Excel function, NORMSDIST(1.2)= .8849 and NORMSDIST(0)= .5